### v1

# HallE-Switch: Rethinking and Controlling Object Existence Hallucinations in Large Vision Language Models for Detailed Caption

Abstract
Current large vision-language models (LVLMs) achieve remarkable progress, yet there remains significant uncertainty regarding their ability to accurately apprehend visual details, that is, in performing detailed captioning. To address this, we introduce CCEval, a GPT-4 assisted evaluation method tailored for detailed captioning. Interestingly, while LVLMs demonstrate minimal object existence hallucination in existing VQA benchmarks, our proposed evaluation reveals continued susceptibility to such hallucinations. In this paper, we make the first attempt to investigate and attribute such hallucinations, including image resolution, the language decoder size, and instruction data amount, quality, granularity. Our findings underscore the unwarranted inference when the language description includes details at a finer object granularity than what the vision module can ground or verify, thus inducing hallucination. To control such hallucinations, we further attribute the reliability of captioning to contextual knowledge (involving only contextually grounded objects) and parametric knowledge (containing inferred objects by the model). Thus, we introduce HallE-Switch, a controllable LVLM in terms of Hallucination in object Existence. HallE-Switch can condition the captioning to shift between (i) exclusively depicting contextual knowledge for grounded objects and (ii) blending it with parametric knowledge to imagine inferred objects. Our method reduces hallucination by 44% compared to LLaVA and maintains the same object coverage.
1 Introduction
In recent years, Large Vision-Language Models (LVLMs) (Liu et al., 2023a; Dai et al., 2023; Li et al., 2023b; Zhu et al., 2023) have achieved significant progress, advancing tasks such as detailed captioning, visual conversations, and vision question-answering (VQA) (Goyal et al., 2017; Liu et al., 2023b; Hudson & Manning, 2019; Fu et al., 2023). However, similar to Large Language Models (LLMs) (Touvron et al., 2023a; Team, 2023; OpenAI, 2022) in the NLP domain, LVLMs confront the issue of hallucination. This challenge is especially severe in detailed image captioning, crucial for accurately describing visual details and supporting applications in robotics (Huang et al., 2023), visual search (Hu et al., 2023), etc. To better understand and address this challenge, we first outline the various types of hallucinations frequently observed in the detailed captions:
Object Existence Hallucination: The detailed image description references objects that are not present.
Object Attribute Hallucination: The detailed image description inaccurately characterizes objects, misrepresenting attributes such as color, shape, and size.
Object Relationship Hallucination: The detailed image description inaccurately depicts the relationships or interactions among objects, including erroneous relative positions, interaction states, and actions involving two or more objects.
Evaluating detailed captions is inherently complex. Some of the efforts, including benchmarks like POPE (Li et al., 2023d), evaluate object hallucination using VQA. Such a bias towards VQA-based evaluations might result in an incomplete assessment of detailed captions, particularly in obtaining a comprehensive view of visual details. To bridge this gap, we introduce CCEval, designed specifically for object existence hallucination in detailed captions. To avoid the model gaining an unfair advantage by favoring shorter descriptions, CCEval maintains consistency in metrics such as average sentence length and the number of objects. Notably, even models that well-performed on VQA-based object hallucination benchmarks showed substantial hallucinations when evaluated with CCEval. Given the finding, we systematically tackle hallucinations, with an initial focus on object existence hallucination.
In our exploration to uncover the underlying causes of object existence hallucination, we find various factors including the size of the language decoder, the quantity, quality, and granularity of instruction data, and the input resolution to the vision encoder. We observe that increasing the language decoder size can reduce hallucination, but more crucially, vision encoder needs to ground the object accurately. As illustrated in Figure 1(a), during training of LVLMs, the goal is to establish a one-to-one correspondence between objects mentioned in the caption and those present in the image. Objects successfully grounded by the vision encoder form accurate associations, internalizing them as contextual knowledge. Conversely, objects in the language that the vision encoder fails to ground create word-word semantic associations, which can be attributed to the generalization from the parametric knowledge within the model’s parameters. During inference, when the model draws from such parametric knowledge, any misalignment can manifest as hallucination, as the model attempts to “guess” details not grounded by the vision module.
Recognizing the significance of both contextual and parametric knowledge in ensuring generation reliability, we present HallE-Switch, a novel approach to control the extent of expressed hallucination or parametric knowledge. The motivation is that not all hallucination is bad, and it is more desirable to control the generalization rather than outright removal of all imagined objects. We curate a 33k dataset similar to LLaVA (Liu et al., 2023a), incorporating both pure contextual knowledge and a blend of contextual knowledge with marked parametric knowledge. Leveraging this dataset, we train a lightweighted single linear layer to control over the frozen LVLM. As demonstrated in Figure 1(b), a singular continuous parameter adjustment (e.g. ) during inference enables the model to produce detailed captions with only contextual knowledge (e.g., ) or blend with parametric knowledge (e.g., ). Furthermore, the inferred objects from parametric knowledge are automatically highlighted with distinct tokens (e.g., [object]) for human reference. This method offers the advantage of preserving object count and coverage as well as sentence length, while effectively control object existence hallucination.
Overall, our contributions are:
A novel evaluation method for detailed caption object existence hallucination, with metrics such as object count, coverage, and average sentence length, alongside hallucination assessment;
A comprehensive analysis on LVLM components that influence hallucination, with a specific focus on alignment issues in the vision encoder and instruction data;
An approach to control object existence hallucination within detailed captions.
2 Hallucination Analysis
Object existence hallucination can be influenced by several factors, including the language decoder, instruction data, and vision encoder. In our analysis, we address each factor individually. For a diverse methodological exploration, we select LLaVA, InstructBLIP (Dai et al., 2023), and Shikra (Chen et al., 2023): LLaVA and Shikra share the same model structure; Shikra and InstructBLIP use mixed-dataset and multi-task instruction data; InstructBLIP finetunes only Q-former, while the other finetune projector and LLM.
LLaVA uses a linear projector to map visual token as a soft-prompt into LLM input tokens. LLaVA has a two-stage training, where the initial stage focuses on simple caption pretraining solely for the linear projector, while the subsequent stage finetunes both the projector and LLM on instruction data. Instruction data leverages language-only GPT-4 by inputting visual ground truth from COCO dataset.
InstructBLIP adopts the BLIP-2 architecture, and is distinguished by its training of a Q-former, which bridges the frozen vision encoder and LLM. InstructBLIP’s instruction fine-tuning spans across 26 distinct datasets.
Shikra mirrors LLaVA’s model structure. It eliminates the pretrain stage, but introduce grounding task during finetuning. Shikra is trained on multiple datasets like InstructBLIP.
2.1 Benchmarks
There are two primary approach, VQA-based and caption-based benchmarks, for evaluating object existence hallucination in LVLMs.
| Benchmark | Model | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | LLaVA | 73.13 | 66.95 | 94.53 | 78.39 | 72.78 |
| LLaVA | 78.49 | 73.57 | 90.93 | 81.34 | 63.71 | |
| Shikra | 86.99 | 94.77 | 79.13 | 86.25 | 43.04 | |
| InstructBLIP | 86.60 | 80.74 | 96.13 | 87.77 | 59.53 | |
| POPE - Popular | LLaVA | 59.87 | 55.88 | 93.80 | 70.03 | 83.93 |
| LLaVA | 70.80 | 64.73 | 91.40 | 75.79 | 70.60 | |
| Shikra | 84.35 | 88.10 | 79.43 | 83.54 | 45.08 | |
| InstructBLIP | 71.27 | 64.20 | 96.13 | 76.99 | 74.87 | |
| POPE - Adversarial | LLaVA | 57.06 | 54.07 | 93.93 | 68.63 | 86.87 |
| LLaVA | 63.93 | 59.03 | 91.07 | 71.63 | 77.13 | |
| Shikra | 82.88 | 85.32 | 79.43 | 82.27 | 46.55 | |
| InstructBLIP | 72.10 | 65.13 | 95.13 | 77.32 | 73.03 | |
| Benchmark | Model | Existence | Count | Position | Color | Total |
| MME | LLaVA | 150.00 | 48.33 | 50.00 | 55.00 | 303.33 |
| LLaVA | 180.00 | 113.33 | 55.00 | 95.00 | 443.33 | |
| Shikra | 185.00 | 118.33 | 75.00 | 155.00 | 533.33 | |
| InstructBLIP | 185.00 | 60.00 | 50.00 | 125.00 | 420.00 |
VQA-based benchmarks pose questions about details of objects within images. For a model to be considered hallucination-free, it should address these visual questions accurately. Notably, a large proportion of these questions are simply binary, typically asking about the presence or attributes of objects.
The POPE benchmark evaluates object existence hallucination by a polling-based query method, consisting of a series of yes/no questions on sampled objects from visual instructions. POPE contains three sets: random, popular, and adversarial. These subsets respectively focus on randomly selected objects, frequently occurring objects, and those objects that co-occur in training sets. We choose POPE evaluation on MSCOCO (Lin et al., 2014) dataset. MME coarse-grained recognition construct yes/no questions similarly but selects objects at random. This benchmark has 30 images, with each image paired with two questions: one positive and one negative.
In Table 1, LLaVA exhibits the greatest degree of hallucination, whereas Shikra outperforms other models in both POPE and MME. Specifically, Shikra shows a significantly higher F1 score in both POPE-popular and POPE-adversarial categories, while LLaVA displays the lowest. Additionally, Shikra’s ”Yes” ratio is closer to a balanced 50% compared to other models. However, in subsequent sections, we demonstrate that these observations from VQA-based benchmarks are not consistent with those from caption-based benchmarks.
| Model | CHAIR | CCEval (Ours) | |||||||
| CHAIR | CHAIR | Avg. Length | Avg. Object | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object | |
| LLaVA | 24.1 | 9.1 | 42.5 | 3.7 | 72.00 | 19.7 | 32.74 | 92.27 | 9.19 |
| LLaVA | 60.6 | 18.4 | 90.2 | 7.6 | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 |
| Shikra | 59.1 | 16.6 | 91.2 | 7.5 | 83.00 | 24.40 | 33.29 | 109.37 | 9.10 |
| InstructBLIP | 1.4 | 1.7 | 2.3 | 0.8 | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 |
Caption-based benchmarks, like CHAIR, begin by splitting the sentence and extracting nouns. Subsequently, it augments the ground truth objects by incorporating hard-coded synonyms and phrases, forming a ground truth set. The benchmark then identifies hallucinated objects by comparing the objects in the caption with this ground truth set. CHAIR computes and as follows:
Table 2(left) reveals that while InstructBLIP exhibits minimal object existence hallucination, it averages a mere 0.8 objects per sentence. In contrast, LLaVA and Shikra manifest a higher degree of hallucination, but they also generate more detailed captions, outputting as many as 7.6 and 7.5 objects per sentence, respectively. We find infeasible to comparing object hallucinations is impractical when there is a significant disparity in average sentence length and the number of objects.
Apart from the disparities in the average number of objects, the use of a hard-coded ground truth set is another challenge. To counter these challenges, we introduce CCEval, a GPT-4 assisted evaluation for detailed captions. We first prompt LVLMs to generate detailed captions. Subsequently, utilizing GPT-4’s in-context learning capabilities, we extract individual objects from these captions and identify hallucinated ones by referencing the provided ground truth objects. On top of CHAIR, we introduce ”coverage” metric to ensure that the captions are detailed enough. This metric computes the ratio of objects in the caption that match the ground truth to the total number of ground truth objects from Visual Genome (Krishna et al., 2017). We additionally record and and balance the average number of objects and the average length of captions across all cases.
As reflected in Table 2(right), when subjected to consistent constraints—average sentence length approximately 100 words and around 9 objects per sentence—all models exhibit comparable results. Interestingly, while Shikra surpass other models in VQA-based benchmarks, especially in the POPE , it under-performed in CCEval. This suggests that object existence hallucination in detailed captions is consistently captured by VQA-based evaluations.
| Benchmark | Model | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | LLaVA | 75.77 | 69.79 | 93.47 | 79.91 | 69.04 |
| LLaVA | 78.49 | 73.57 | 90.93 | 81.34 | 63.71 | |
| LLaVA | 78.14 | 73.18 | 90.93 | 81.09 | 64.05 | |
| InstructBLIP | 86.60 | 80.74 | 96.13 | 87.77 | 59.53 | |
| InstructBLIP | 88.73 | 86.67 | 92.33 | 89.41 | 54.91 | |
| POPE - Popular | LLaVA | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 |
| LLaVA | 70.80 | 64.73 | 91.40 | 75.79 | 70.60 | |
| LLaVA | 72.43 | 66.45 | 90.60 | 76.67 | 68.17 | |
| InstructBLIP | 71.27 | 64.20 | 96.13 | 76.99 | 74.87 | |
| InstructBLIP | 80.53 | 74.70 | 92.33 | 82.59 | 61.80 | |
| POPE - Adversarial | LLaVA | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
| LLaVA | 63.93 | 59.03 | 91.07 | 71.63 | 77.13 | |
| LLaVA | 66.30 | 60.91 | 91.00 | 72.98 | 74.70 | |
| InstructBLIP | 72.10 | 65.13 | 95.13 | 77.32 | 73.03 | |
| InstructBLIP | 73.97 | 67.53 | 92.33 | 78.01 | 68.37 | |
| Benchmark | Model | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | LLaVA | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
| LLaVA | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 | |
| LLaVA | 82.00 | 21.80 | 31.26 | 106.85 | 9.07 | |
| InstructBLIP | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 | |
| InstructBLIP | 64.00 | 16.70 | 33.60 | 101.63 | 8.06 |
2.2 Language Decoder
We investigate if expanding the size of the language backbone can mitigate object existence hallucination. As detailed in Table 3, the language decoder of LLaVA is increased from 7B to 33B, and for InstructBLIP, it is increased from 7B to 13B. The result shows that hallucination for LLaVA reduced for POPE but not for CCEval. For InstructBLIP, and is reduced by 8 and 5.6 on CCEval, respectively. More importantly, although there is a gain for scaling up language backbone, the performance does not scale linearly, as the accuracy increased 5.73 from 7B to 13B and only 2.37 from 13B to 33B on POPE-adversarial.
2.3 Data
Similar to our approach with the language decoder, we begin by scaling up the volume of instruction finetuning data, ranging from 80K to 2.4M. As illustrated in Table 4, the LLaVA model, finetuned on 80K instruction data, exhibits fewer object existence hallucinations compared to the models finetuned on 150K and SVIT (Zhao et al., 2023). The result suggests extra data may increase hallucination for both VQA-based and caption-based evaluations. Given that all three datasets are generated by GPT-4, we question the quality of the data. LRV also raises this concern, suggesting that the training data itself might contain hallucinations. Some examples of training data are presented in the Appendix. Interestingly, our examination shows no object existence hallucination: the objects in the captions are contained the ground truth objects from MSCOCO. However, we identify certain ground truth objects are challenging for human observers to ground, due to factors like size, resolution, and occlusion. This led us to hypothesize that the vision encoder might also struggle to ground these objects effectively.
| Benchmark | Finetune Data | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | 80K | 73.13 | 66.95 | 94.53 | 78.39 | 72.78 |
| 158K | 75.77 | 69.79 | 93.47 | 79.91 | 69.04 | |
| SVIT | 52.34 | 52.00 | 97.87 | 67.92 | 97.01 | |
| POPE - Popular | 80K | 59.87 | 55.88 | 93.80 | 70.03 | 83.93 |
| 158K | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 | |
| SVIT | 50.77 | 50.43 | 90.47 | 64.76 | 89.70 | |
| POPE - Adversarial | 80K | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
| 158K | 58.47 | 55.00 | 93.07 | 69.14 | 84.6 | |
| SVIT | 51.37 | 50.77 | 90.33 | 65.00 | 88.97 | |
| Benchmark | Model | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | 80K | 72.00 | 19.70 | 32.74 | 92.27 | 9.19 |
| 158K | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 | |
| SVIT | 87.00 | 23.30 | 47.46 | 296.63 | 18.14 |
| Benchmark | Vision Encoder | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | CLIP-L-112x | 79.00 | 21.70 | 32.04 | 110.36 | 9.12 |
| CLIP-L-224x | 74.00 | 19.30 | 32.83 | 113.03 | 9.18 | |
| CLIP-L-336x | 64.00 | 16.00 | 33.37 | 108.52 | 8.92 |
| Benchmark | Vision Encoder | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | CLIP-L-224x | 79.00 | 21.70 | 32.04 | 110.36 | 9.12 |
| CLIP-L-336x | 79.00 | 18.90 | 36.00 | 111.55 | 9.19 | |
| CLIP-L-224x (SW) | 72.00 | 18.70 | 36.89 | 110.43 | 8.65 |
2.4 Vision Encoder
Intuitively, increasing image resolution enhances model’s perception of finer details, thus making the grounding of objects mentioned in the caption easier. To verify our hypothesis from the previous section, we increase the input image resolution for the vision encoder. Specifically, for our evaluation, the resolution for LLaVA was incremented from 224x to full resolution using a sliding window approach for efficiency, as detailed in the Appendix. Table 6 shows a constant decrease in hallucination and increase in object coverage. Additionally, we assess LLaVA with Llama 2, varying the resolution from 112x to 336x. For the 112x112 resolution, the original image was downscaled to 112x and subsequently upscaled to 224x before being input to CLIP-Large-224x. Table 5 gives a consistent observation that larger input resolution can reduce hallucination.
Conclusion. Through our systematic analysis of object existence hallucination, we summarize several insights: (1) Enlarging the language decoder does mitigate hallucination, but gains diminish while scaling up, indicating it is not the primary factor. (2) Expanding the volume of instruction data actually increases hallucination. Upon inspection of training data, we find certain objects described in the captions might not be grounded by the vision encoder. (3) Improving input image resolution significantly reduces hallucination by enhancing grounding, validating our hypothesis.
Reflecting on these findings, we attempt to provide an explanation for the reason of object existence hallucination in detailed captions. The process of image captioning in LVLMs can be perceived as a form of information mapping or translation. Ideally, the goal is to have a direct one-to-one correspondence between objects identified in the image and those mentioned in the captions. Objects successfully grounded by the vision encoder form accurate associations. We define this kind of knowledge as contextual knowledge in the model, following (Neeman et al., 2023). We define parametric knowledge, the knowledge is encoded in the model’s parameter, to be the association of objects in the language with other words instead of with corresponding image object feature, formed when they fails to ground by the vision encoder. During inference, when the model draws from this parametric knowledge, it attempts to “guess” details not grounded by the vision module and is perceived as object existence hallucination. However, such guessing is necessary depending the details required by the downstream task. Therefore, our approach work towards controlling and balancing parametric knowledge rather than eliminating.
3 Hallucination Controlling
We introduce HallE-Switch, a LVLM designed to control the extent of parametric knowledge within detailed captions. For this purpose, we developed two datasets: the first captures solely contextual knowledge, while the second merges both contextual and parametric knowledge. Using these datasets, we integrated a control projector into the model to control parametric knowledge.
3.1 Data Generation
We present an approach to generating new data on MSCOCO (Sharma et al., 2018) source labels using GPT-4, follwing LLaVA’s methodology.
Grouping using RAM. We begin by passing MSCOCO’s ground truth objects utilizing the open vocabulary detector RAM (Zhang et al., 2023). This step aims to simulate the maximum visual granularity achievable by a vision encoder. RAM categorizes the data into sets: “grounded” (contextual group) and “not grounded” (parametric group).
Contextual Data Generation. For our first dataset, we adhere to LLaVA’s original detailed caption creation pipeline. This involves generating detailed captions using only objects from the contextual group. To do this, we feed MSCOCO source labels (including object classes, bounding boxes, and short captions) into GPT-4.
Parametric Joint Data Generation. The second dataset incorporates both contextual and parametric knowledge. Here, we begin with LLaVA’s original detailed captions and annotate objects from the parametric group with special tokens. Specifically, we enclose the ’undetected’ objects with brackets. Formally, if denotes the original image caption sentence and represents a set of non-grounded objects, our data processing can be represented as:
The purpose of bracketing the parametric objects is twofold: it serves as an indicator during inference and provides a hint during training.
3.2 Hallucination Switch
Inspired by LM-Switch (Han et al., 2023), we propose to address hallucination by learning a control parameter that serves as a “switching value”, e.g., for permitted imagination and for restricting imagination, as depicted in Figure 2(a). Let represent the LLM with fixed parameters. Its output word embedding is modified as , leading to the derived model as:
The learned can be regarded as the transformation from a generic word space to the object sensitive word space, where word-word semantic correspondence is optimized to object correspondence, and governs the intensity of imagining related objects.
Training. To train such a controlling parameter, we leverage the contrastive training data covering both contextual and parametric datasets in Sec 3.1. For data with only contextual knowledge, we assign when inputting into the model. In contrast, for data with both contextual and parametric knowledge, we use . Notably, only the linear layer is fine-tuning throughout the training phase.
Inference. At the inference stage, can adopt any value within the interval . Specifically, an value of -1 corresponds to minimal reliance on parametric knowledge, whereas a value of 1 indicates a strong inclination towards such knowledge. Detailed theoretical explanation are elaborated upon in the Appendix.
4 Experiment
| Setting | LLM | Resolution | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| 158K baseline | LLaVA | 224x | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
| only ind. | LLaVA | 224x | 53.00 | 63.90 | 12.01 | – | 1.66 |
| w/o ind. | LLaVA | 224x | 57.00 | 17.10 | 37.60 | 108.94 | 7.63 |
| w/ ind. | LLaVA | 224x | 57.00 | 14.00 | 37.62 | 108.94 | 9.22 |
| 158K baseline | LLaVA Llama 2 | 336x | 64.00 | 16.00 | 33.37 | 108.52 | 8.92 |
| only ind. | LLaVA Llama 2 | 336x | 52.00 | 62.31 | 19.90 | – | 1.3 |
| w/o ind. | LLaVA Llama 2 | 336x | 52.00 | 11.62 | 34.70 | 106.94 | 7.23 |
| w/ ind. | LLaVA Llama 2 | 336x | 52.00 | 9.86 | 39.31 | 106.94 | 8.52 |
| Switch | LLM | Resolution | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| 1 | LLaVA | 224x | 89.00 | 26.60 | 32.80 | 108.54 | 9.72 |
| 0.5 | LLaVA | 224x | 85.00 | 27.92 | 34.02 | 109.33 | 8.81 |
| -0.5 | LLaVA | 224x | 81.00 | 24.88 | 35.87 | 118.08 | 8.04 |
| -1 | LLaVA | 224x | 76.00 | 20.90 | 33.88 | 133.79 | 8.02 |
| 1 | LLaVA Llama 2 | 336x | 65.00 | 14.58 | 36.14 | 102.18 | 8.37 |
| 0.5 | LLaVA Llama 2 | 336x | 65.00 | 14.44 | 32.32 | 103.51 | 8.45 |
| -0.5 | LLaVA Llama 2 | 336x | 66.00 | 13.79 | 33.07 | 105.57 | 8.41 |
| -1 | LLaVA Llama 2 | 336x | 43.00 | 6.37 | 34.37 | 136.28 | 8.79 |
4.1 Finetune on Parametric Joint Data
Before we present experiments on HallE-Switch, we show the upper-bound experiment result on how well the model can indicate parametric knowledge. We directly finetune the LLaVA model on parametric joint data. Intuitively, the model is trained on data indicating parametric knowledge. Its output should identify parametric knowledge accurately. Specifically, the model should put a bracket around every ”guessed” objects for indication of hallucination.
Therefore, we evaluate the object hallucination in three different settings: 1. Evaluation only on indicated objects: We do CCEval which evaluate CHAIR, CHAIR, and coverage on objects inside the bracket. The result should reflect a high level of hallucination. 2. Evaluation without indicated objects: We do CCEval which evaluate CHAIR, CHAIR, and coverage on objects not in the bracket. The result should reflect a low level of hallucination. 3. Evaluation with indicated objects. We do CCEval as following:
Evaluation only on indicated objects. The hallucination level for indicated objects, denoted as , is 63.90 for LLaVA and 62.31 for LLaVA. It is considerably higher than the baselines all other models. Concurrently, their coverage is 12.01 and 19.09 for LLaVA and LLaVA, respectively, which both of them significantly lower than the coverage of 33.58 for LLaVA baseline model. The experiment results indicate the object within special tokens has significantly higher hallucination rate and lower coverage rate which support our assumption that the object inside the indication tokens are objects indicated from parametric knowledge.
Evaluation without indicated objects. For objects outside of the special token scope, we found that hallucination is markedly reduced, CHAIR decreased from 82 to 57 which is 30.5% percent improvement and CHAIR decrease 32.4%, from 25.3 to 17.10, compared to the baseline. This suggests that the model is less prone to make erroneous assumptions for objects not marked by brackets.
Evaluation with indicated objects.: We observe a significant decline in the hallucination without any reduce in object coverage. LLaVA CHAIR improved from 25.30 to 14.00 which has 44.66% improvements. For LLaVA CHAIR improved from 16 to 9.86 which also has 38.38% improvements.
4.2 Hallucination Controlling
During model inference, we select 4 different , ranging from to . As shown in Table 8, we evaluate HallE-Switch and HallE-Switch model, which use LLaVA as backbone. For the 7B model, we train it by removing the special token in parametric joint data, showing that indication is not a necessary condition for switch to work. means the switch trained purely on contextual only data, which try to minimize the hallucination, where switch to maximize parametric knowledge output. The results show that as increase, CHAIR increase from 20.90 to 26.6, the coverage keeps at a similar level.
For the 13B model, we keep the indication inside the parametric joint data. The HallE-Switch achieves the best in object existence hallucination metric. With switch set to -1 and indication, we have versus baseline’s 64 and vs baseline’s 16, and the coverage of the model is not decreasing.
5 Related Work
Large Vision-Language Models (LVLMs).
The rapid advancements in Large Language Models (LLMs) (Touvron et al., 2023a; Chung et al., 2022; Touvron et al., 2023b; Anil et al., 2023; Driess et al., 2023; Scao et al., 2022) combined with a surge in open-source initiatives, such as GPTs, LLaMAs, MPT, and Vicuna, have paved the way for the emergence of extensive vision-language models Liu et al. (2023a); Goyal et al. (2017); Li et al. (2023b); Zhu et al. (2023); Sun et al. (2023); Ye et al. (2023); Bai et al. (2023); Chen et al. (2023); Peng et al. (2023) like GPT-4 OpenAI (2023). LLaVA introduced the concept of integrating a simple projector during LLM fine-tuning. Concurrently, BLIP2 (Li et al., 2023c)presented q-former-based LVLMs, with InstructBLIP (Dai et al., 2023) further refining the concept. Otter aims to improve OpenFlamingo’s Alayrac et al. ; Awadalla et al. (2023) directive adherence. MiniGPT-4 Zhu et al. (2023) highlights the potential of sophisticated LLMs and advocates for a unified layer to synchronize visual and text models, offering GPT4-like capabilities but with reduced computational demands. mPLUG-Owl Ye et al. (2023) suggests a two-step method: first aligning visual aspects, then refining the language model using techniques like LoRA. QWen-VL Bai et al. (2023) scales up LMM pre-training significantly and showcases stellar performance on various benchmarks.
Hallucination
Before the emergence of LLMs, the NLP community characterized “hallucination” as the generation of content that either lacked coherence or deviated from its original source (Ji et al., 2023a). Hallucinations have been widely studied in traditional NLG (Ji et al., 2023b) tasks, including machine translation (Zhou et al., 2020; Lee et al., 2019), data-to-text (Rebuffel et al., 2021; Kasner & Dušek, 2022; Lee et al., 2022), summarization (Cao et al., 2022), dialogue (Dziri et al., 2022) and QA (Shuster et al., 2021). After LLMs significantly bring model performance to next stage, several works bring new meaning to hallucination as the generation of text that is erroneous, nonsensical, or detached from reality (Shi et al., 2023; Lin et al., 2021). For LVLMs, previous studies have been mainly focusing on object hallucination (Marino et al., 2019; MacLeod et al., 2017; Li et al., 2023a; d).
6 Conclusion
In summary, this study delves deep into the object hallucination phenomena within the detailed captions of LVLMs, advancing understanding of the accuracy and unwarranted inference in describing visual details. We introduce a novel and comprehensive evaluation method for object existence hallucination in detailed captions. We conduct an in-depth and component-wise analysis of LVLMs, meticulously examining each element that might result in hallucination. We further identify an alignment issue between the vision encoder and the instruction data. To alleviate such hallucination, we introduce controlling parameters over LVLMs to condition the inference of objects.
Appendix A Appendix
A.1 Training Data Quality
We sampled three images from the MSCOCO dataset, as illustrated in Table 9. For each image, we present the visual content, a detailed caption generated by GPT-4 based on bounding boxes and regional captions, and the object ground truth labels derived from the MSCOCO dataset annotations.
First Image. This image showcases three remote controls, posing a unique challenge. The ambiguity lies in distinguishing the type of remote, be it for video games or televisions. Additionally, the remotes near the wall are relatively small, making them harder to see. A person is partially visible in the image, with only their knees being evident. There is also an incomplete bottle on the image’s left side.
Second Image. A significant issue with this image is the individuals visible behind a window. Not only detection models struggle to recognize them, even for human observers, counting the individuals inside the train is challenging. This particular issue is prevalent in many of MSCOCO’s traffic-related images.
Third Image. This image depicts a table around which two individuals are seated. Close observation is required to recognize both individuals. A clear indication of one person is a pair of hands, while the other individual is considerably harder to spot. Additionally, there seems to be an annotation error in the ground truth labels: it indicates only one bowl, neglecting to include plates and other items. Upon closer inspection, there are four plates.
Image |
Caption |
GT Labels |
The image depicts a cozy entertainment room featuring a flat screen TV situated against the wall. A dining table can be seen next to the TV, with some bottles and a cup placed on top of it. Multiple video game controllers, including Wii remotes, are also present on the dining table, suggesting that people are using the TV to play video games. To the far right of the scene, another remote is located close to the wall. A person appears to be sitting or standing behind the dining table, likely either watching the game or waiting for their turn to play. The assortment of bottles on the table suggests that the guests are enjoying drinks during their video game session. |
tv: 1 cup: 1 remote: 3 bottle: 4 person: 1 dining table: 1 |
In the image, an orange mass transit trolley is making its way through a city. A person is crossing the street in front of the trolley while holding a garbage bag, appearing to be cautious about the approaching vehicle. Another person is standing close to the person crossing the street, and there are two more individuals nearby. In the scene, various vehicles surround the trolley, including cars, buses, and trucks. One of the cars is parked right behind the trolley, while another is situated farther back. Two buses can be seen, with one staying behind the trolley and the other on the right side of it. A truck is also present at the far left side of the scene. A woman nearby is holding a handbag, completing the busy urban setting. |
car: 3 bus: 2 truck: 1 train: 1 person: 5 handbag: 1 |
The image shows a table in a restaurant, which appears to have been recently used for a meal. The table is set with four place settings, including white dishes. On the table, there are multiple cups, a bowl, and a mix of silverware like forks and spoons laid out. Some of the dishes look dirty, with used napkins and eating utensils scattered around. In the background, two people can be seen, one situated on the left side and another on the right side behind the table. There are also chairs located near the table, with one chair positioned close to the left side and another chair closer to the right side of the frame. |
cup: 5 spoon: 3 fork: 2 person: 2 chair: 2 bowl: 1 dining table: 1 |
A.2 Prompt
Table 10 is the prompt we use for caption object extraction.
Table 11 is the prompt we use for hallucination object matching.
Table 12 is the prompt we use for finding ground truth object coverage.
A.3 Sliding Window
We pad the original image to dimensions of . Then, we divide the image into a grid, where each cell measures . The encoder processes these cells sequentially, starting from the top-left and moving towards the bottom-right. The visual tokens from each cell are then concatenated.
A.4 Experiment Settings
The LLaVA we used in all experiments is pretrained on on LCS-558k which is subset of LAION (Schuhmann et al., 2021), CC (Sharma et al., 2018) and SBU (Ordonez et al., 2011) data, and finetuned on Instruct-158K instruction data. We use Vicuna version 1.3 as initialized language decoder and CLIP-Large as vision encoder. The 158K finetuning data consists of detailed caption, complex reasoning, and conversation.
We employed the RAM detector, specifically the RAM-14M variant, which uses a Swin-Large backbone. In the data generation stage, we focused on the ’detailed caption 23K’ file from the LLaVA Instruction set comprising 158K entries. This file was generated using a specific prompt provided by LLaVA repository111https://github.com/haotian-liu/LLaVA/tree/main/playground/data/prompts/detail_description. Our preprocessing involved adding brackets ’[]’ around objects filtered by the RAM detector to uniquely identify them in the captions.
Hallucination switch only involves only finetuning a linear layer added before layer. We freeze all other layers and only finetune switch layer with 3 epochs. The dataset is generated data with associated value. We have 10K contextual only detailed caption data and 23K parametric joint detailed caption data.
A.5 Theoretical Explanations
HallE-Switch Formulation
HallE-Switch follow LLaVA’s training strategy, which freezes vision encoder and language model, only finetune the projector for the first stage. The visual encoder transfer image to visual features:
The projector connects image feature to the word embedding space:
In the second stage, both projector and language model is trained. The model is trained on two type of data: 1. Contextual only data. 2. Contextual and parametric combined data.
Language model trained on contextual only data has a distribution initiating from . The other language model trained on combined data has a distribution initiating from
With the Theorem 1 in LM-Switch (Han et al., 2023), under the same assumptions, there exists an matrix , transforming a word embedding to , which is equivalent to let a LM simulate the text distribution initiating from another distribution.
Inspired by the Theorem, we propose a linear transform in word embedding space for LVLM. Let be the finetuned LLM, with switch layer/projector denoted as , we replace each word embedding with , making the new language model as HallE-Switch’s language decoder. is adjustable from -1 to 1. We assign and fit contextual only data; assign with the contextual and parametric combined data. After finetuning the HallE-Switch, the user only needs to specify a switch value and do normal vision language task like image captions. We use maximal likelihood as the training objective.
Continuous Control
The design of LM-Switch maintains a linearity guarantee, with proof of the switch model’s distribution is close to a linear interpolation. Let be the maximum eigen-value of W. When varying s value,
distribution of the switch model is close a linear interpolation of and , meaning that the model distribution changes linearly. Therefore, our method can take any between 1 and -1.
![[Uncaptioned image]](extracted/5143148/figures/000000146988.jpg)
![[Uncaptioned image]](extracted/5143148/figures/000000157923.jpg)
![[Uncaptioned image]](extracted/5143148/figures/000000192920.jpg)