### v3

# HallE-Control: Controlling Object Hallucination in Large Multimodal Models

Abstract
Current Large Multimodal Models (LMMs) achieve remarkable progress, yet there remains significant uncertainty regarding their ability to accurately apprehend visual details, that is, in performing detailed captioning. To address this, we introduce CCEval, a GPT-4 assisted evaluation method for detailed captioning. Interestingly, while LMMs demonstrate minimal object existence hallucination in existing VQA benchmarks, our proposed evaluation reveals continued susceptibility to such hallucinations. In this paper, we make the first attempt to investigate such hallucination from different aspects, including image resolution, the language decoder size, and instruction data amount, quality, granularity. Our findings underscore the unwarranted inference when the language description includes details at a finer object granularity than what the vision module can ground or verify, thus inducing hallucination. To control such hallucinations, we further attribute the reliability of captioning to contextual knowledge (involving only contextually grounded objects) and parametric knowledge (containing inferred objects by the model). Thus, we introduce HallE-Control, a controllable LMM in terms of Hallucination in object Existence. HallE-Control can condition the captioning to shift between (i) exclusively depicting contextual knowledge for grounded objects and (ii) blending it with parametric knowledge to imagine inferred objects. Our method reduces hallucination by 44% compared to LLaVA7B and maintains the object coverage. Our code is publicly available at https://github.com/bronyayang/HallE_Control
1 Introduction
In recent years, Large Multimodal Models (LMMs) (Liu et al., 2023d; Dai et al., 2023; Li et al., 2023b; Zhu et al., 2023) have achieved significant progress, advancing tasks such as detailed captioning, visual conversations, and vision question-answering (VQA) (Goyal et al., 2017; Liu et al., 2023e; Hudson & Manning, 2019; Fu et al., 2023). However, similar to Large Language Models (LLMs) (Touvron et al., 2023a; Team, 2023; OpenAI, 2022) in the NLP domain, LMMs confront the issue of hallucination. This is particularly severe in detailed image captioning, which hinders the performance of downstream applications in robotics (Huang et al., 2023), visual search (Hu et al., 2023), etc. To better understand and address this challenge, we first outline three types of object hallucinations frequently observed in the detailed captions: (1) Object Existence Hallucination - The detailed image description references objects that are not present; (2) Object Attribute Hallucination - The detailed image description inaccurately characterizes objects, misrepresenting attributes such as color, shape, and size; (3) Object Relationship Hallucination - The detailed image description inaccurately depicts the relationships or interactions among objects, including erroneous relative positions, interaction states, and actions involving two or more objects. In this work, we mainly focus on defining the metric, analyzing the cause, and addressing the the problem of object existence hallucination.
Evaluating detailed captions is inherently complex. Some of the efforts, including benchmarks like POPE (Li et al., 2023e), evaluate object hallucination using VQA. Such a bias towards VQA-based evaluations might result in an incomplete assessment of detailed captions, which requires obtaining a comprehensive view of visual details. To bridge this gap, we introduce CCEval, designed specifically for object existence hallucination in detailed captions. To avoid the model gaining an unfair advantage by favoring shorter descriptions, CCEval maintains consistency in metrics such as average sentence length and the number of objects (see Sec. 2). Notably, even models that well-performed on VQA-based object hallucination benchmarks showed substantial hallucinations when evaluated with CCEval.
In our exploration to uncover the underlying cause of object existence hallucination, we look into various factors including the size of the language decoder, the quantity, quality, and granularity of instruction data, and the input resolution to the vision encoder. We conclude the most crucial factor to be the alignment between objects mentioned in training caption and those vision encoder can perceive. During training of LMMs, the goal is to establish a one-to-one correspondence between objects mentioned in the caption and those present in the image. Objects successfully grounded by the vision encoder form accurate associations, internalizing them as contextual knowledge. Conversely, objects in the language that the vision encoder fails to ground create word-word semantic associations, which can be attributed to the generalization from the parametric knowledge within the model’s parameters. During inference, when the model draws from such parametric knowledge, any misalignment can manifest as hallucination, as the model attempts to “guess” details not grounded by the vision module.
To address such hallucination, we are motivated by that not all hallucination is bad, and it is more desirable to control the generalization rather than outright removal of all imagined objects. Recognizing the significance of both contextual and parametric knowledge in ensuring generation reliability, we present HallE-Control, a novel approach to control the extent of expressed hallucination or parametric knowledge. We curate a 33k dataset similar to LLaVA (Liu et al., 2023d), incorporating both pure contextual knowledge and a blend of contextual knowledge with marked parametric knowledge. Leveraging this dataset, we train a lightweighted single linear layer to control over the frozen LMM. As demonstrated in Figure 1, a singular continuous parameter adjustment (e.g. ) during inference enables the model to produce detailed captions with only contextual knowledge (e.g., ) or blend with parametric knowledge (e.g., ). Furthermore, the inferred objects from parametric knowledge are automatically highlighted with distinct tokens (e.g., [object]) for human reference. This method offers the advantage of preserving object count and coverage as well as sentence length, while effectively control object existence hallucination.
Overall, our contributions are:
A comprehensive analysis on LMM components that influence hallucination, with a specific focus on alignment issues in the vision encoder and instruction data;
The first approach to control object existence hallucination within detailed captions;
A novel evaluation method for detailed caption object existence hallucination, with metrics such as object count, coverage, and average sentence length, alongside hallucination assessment.
2 Hallucination Analysis
Object existence hallucination can be influenced by several factors, including the language decoder, instruction data, and vision encoder. In our analysis, we address each factor individually. For a diverse methodological exploration, we select LLaVA, InstructBLIP (Dai et al., 2023), and Shikra (Chen et al., 2023): LLaVA and Shikra share the same model structure; Shikra and InstructBLIP use mixed-dataset and multi-task instruction data; InstructBLIP finetunes only Q-former, while the other finetune projector and LLM. More details about models can be found in Appendix.
2.1 Benchmarks
There are two primary methods, VQA-based and caption-based benchmarks, for evaluating object existence hallucination in LMMs.
| Benchmark | Model | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | LLaVA | 73.13 | 66.95 | 94.53 | 78.39 | 72.78 |
| LLaVA | 78.49 | 73.57 | 90.93 | 81.34 | 63.71 | |
| Shikra | 86.99 | 94.77 | 79.13 | 86.25 | 43.04 | |
| InstructBLIP | 86.60 | 80.74 | 96.13 | 87.77 | 59.53 | |
| POPE - Popular | LLaVA | 59.87 | 55.88 | 93.80 | 70.03 | 83.93 |
| LLaVA | 70.80 | 64.73 | 91.40 | 75.79 | 70.60 | |
| Shikra | 84.35 | 88.10 | 79.43 | 83.54 | 45.08 | |
| InstructBLIP | 71.27 | 64.20 | 96.13 | 76.99 | 74.87 | |
| POPE - Adversarial | LLaVA | 57.06 | 54.07 | 93.93 | 68.63 | 86.87 |
| LLaVA | 63.93 | 59.03 | 91.07 | 71.63 | 77.13 | |
| Shikra | 82.88 | 85.32 | 79.43 | 82.27 | 46.55 | |
| InstructBLIP | 72.10 | 65.13 | 95.13 | 77.32 | 73.03 | |
| Benchmark | Model | Existence | Count | Position | Color | Total |
| MME | LLaVA | 150.00 | 48.33 | 50.00 | 55.00 | 303.33 |
| LLaVA | 180.00 | 113.33 | 55.00 | 95.00 | 443.33 | |
| Shikra | 185.00 | 118.33 | 75.00 | 155.00 | 533.33 | |
| InstructBLIP | 185.00 | 60.00 | 50.00 | 125.00 | 420.00 |
VQA-based benchmarks pose questions about objects within images. For a model to be considered hallucination-free, it should address these visual questions accurately. Notably, a large proportion of questions are simply binary, typically asking about the presence or attributes of objects.
The POPE benchmark evaluates object existence hallucination by a polling-based query method, consisting of a series of yes/no questions on sampled objects from visual instructions. POPE contains three sets: random, popular, and adversarial. These subsets respectively focus on randomly selected objects, frequently occurring objects, and those objects that co-occur in training sets. We choose POPE evaluation on MSCOCO (Lin et al., 2014) dataset. MME (Fu et al., 2023) coarse-grained recognition construct yes/no questions similarly but selects objects at random. This benchmark has 30 images, with each image paired with two questions: one positive and one negative.
In Table 1, LLaVA7B exhibits the greatest degree of hallucination, whereas Shikra outperforms other models in both POPE and MME. Specifically, Shikra shows a significantly higher F1 score in both POPE-popular and POPE-adversarial categories, while LLaVA7B displays the lowest. Additionally, Shikra’s ”Yes” ratio is closer to a balanced 50% compared to other models. However, in subsequent sections, we demonstrate that these observations from VQA-based benchmarks are not consistent with those from caption-based benchmarks.
| Model | CHAIR | CCEval (Ours) | |||||||
| CHAIRs | CHAIRi | Avg. Length | Avg. Object | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object | |
| LLaVA7B | 24.1 | 9.1 | 42.5 | 3.7 | 72.00 | 19.7 | 32.74 | 92.27 | 9.19 |
| LLaVA13B | 60.6 | 18.4 | 90.2 | 7.6 | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 |
| Shikra7B | 59.1 | 16.6 | 91.2 | 7.5 | 83.00 | 24.40 | 33.29 | 109.37 | 9.10 |
| InstructBLIP7B | 1.4 | 1.7 | 2.3 | 0.8 | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 |
Caption-based benchmarks, like CHAIR, begin by splitting the sentence and extracting nouns. Subsequently, it augments the ground truth objects by incorporating hard-coded synonyms and phrases, forming a ground truth set. The benchmark then identifies hallucinated objects by comparing the objects in the caption with this ground truth set. CHAIR computes and as follows:
Table 2(left) reveals that while InstructBLIP exhibits minimal object existence hallucination, it averages a mere 0.8 objects per sentence. In contrast, LLaVA13B and Shikra manifest more hallucination, but they also generate more detailed captions, with as many as 7.6 and 7.5 objects per sentence, respectively. We find comparing object hallucinations is impractical when there is a significant disparity in average sentence length and number of objects.
Apart from these disparities, the use of a hard-coded ground truth set is another challenge. To counter these challenges, we introduce CCEval, a GPT-4 assisted evaluation for detailed captions. We first prompt LMMs to generate detailed captions on 100 randomly sampled images from Visual Genome (Krishna et al., 2017). Subsequently, utilizing GPT-4’s in-context learning capabilities, we extract individual objects from these captions and identify hallucinated ones by referencing the provided ground truth objects. On top of CHAIR Rohrbach et al. (2018), we introduce ”coverage” metric to ensure that the captions are detailed enough. This metric computes the ratio of objects in the caption that match the ground truth to the total number of ground truth objects. We additionally record and balance the average number of objects as well as the average length of captions across all cases. More details of CCEval can be found in Appendix.
As reflected in Table 2(right), when subjected to consistent constraints—average sentence length approximately 100 words and around 9 objects per sentence—all models exhibit comparably sub-optimal results. Interestingly, while Shikra surpass other models in VQA-based benchmarks, especially in the POPE , it under-performs in CCEval. This suggests that hallucination in detailed captions is not consistently captured by VQA-based evaluations.
| Benchmark | Model | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | LLaVA | 75.77 | 69.79 | 93.47 | 79.91 | 69.04 |
| LLaVA | 78.49 | 73.57 | 90.93 | 81.34 | 63.71 | |
| LLaVA | 78.14 | 73.18 | 90.93 | 81.09 | 64.05 | |
| InstructBLIP | 86.60 | 80.74 | 96.13 | 87.77 | 59.53 | |
| InstructBLIP | 88.73 | 86.67 | 92.33 | 89.41 | 54.91 | |
| POPE - Popular | LLaVA7B | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 |
| LLaVA | 70.80 | 64.73 | 91.40 | 75.79 | 70.60 | |
| LLaVA | 72.43 | 66.45 | 90.60 | 76.67 | 68.17 | |
| InstructBLIP | 71.27 | 64.20 | 96.13 | 76.99 | 74.87 | |
| InstructBLIP | 80.53 | 74.70 | 92.33 | 82.59 | 61.80 | |
| POPE - Adversarial | LLaVA7B | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
| LLaVA | 63.93 | 59.03 | 91.07 | 71.63 | 77.13 | |
| LLaVA | 66.30 | 60.91 | 91.00 | 72.98 | 74.70 | |
| InstructBLIP | 72.10 | 65.13 | 95.13 | 77.32 | 73.03 | |
| InstructBLIP | 73.97 | 67.53 | 92.33 | 78.01 | 68.37 | |
| Benchmark | Model | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | LLaVA7B | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
| LLaVA | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 | |
| LLaVA | 82.00 | 21.80 | 31.26 | 106.85 | 9.07 | |
| InstructBLIP | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 | |
| InstructBLIP | 64.00 | 16.70 | 33.60 | 101.63 | 8.06 |
2.2 Language Decoder
We investigate if expanding the size of the language backbone can mitigate object existence hallucination. As detailed in Table 3, the language decoder of LLaVA is increased from 7B to 33B, and for InstructBLIP, it is increased from 7B to 13B. The result shows that hallucination for LLaVA reduced for POPE but not for CCEval. For InstructBLIP, and is reduced by 8 and 5.6 on CCEval, respectively. However, although there is a gain for scaling up language backbone, it is not consistent or salient from the observation, suggesting language decoder is not a primary factor in reducing hallucination.
2.3 Data
Similar to our approach with the language decoder, we begin by scaling up the volume of instruction finetuning data, ranging from 80K to 2.4M. As illustrated in Table 4, the LLaVA7B model, finetuned on 80K instruction data, exhibits fewer object existence hallucinations compared to the models finetuned on 150K and SVIT (Zhao et al., 2023a). The result suggests extra data without quality guarantee may increase hallucination for both VQA-based and caption-based evaluations. Given that all three datasets are generated by GPT-4, we question the quality of the data. LRV also raises this concern, suggesting that the training data itself might contain hallucinations. Some examples of training data are presented in the Appendix. Interestingly, we find that GPT-4 does not introduce additional object existence hallucination at all – the issue lies in MSCOCO ground truth objects that are given to GPT-4 and to be asked to strictly include in generated captions. These GT objects are challenging even for human observers to ground, due to factors like size, resolution, and occlusion. This led us to hypothesize that the vision encoder might also struggle to ground these objects effectively.
| Benchmark | Finetune Data | Accuracy | Precision | Recall | F1 | Yes (%) |
| POPE - Random | 80K | 73.13 | 66.95 | 94.53 | 78.39 | 72.78 |
| 158K | 75.77 | 69.79 | 93.47 | 79.91 | 69.04 | |
| SVIT | 52.34 | 52.00 | 97.87 | 67.92 | 97.01 | |
| POPE - Popular | 80K | 59.87 | 55.88 | 93.80 | 70.03 | 83.93 |
| 158K | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 | |
| SVIT | 50.77 | 50.43 | 90.47 | 64.76 | 89.70 | |
| POPE - Adversarial | 80K | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
| 158K | 58.47 | 55.00 | 93.07 | 69.14 | 84.6 | |
| SVIT | 51.37 | 50.77 | 90.33 | 65.00 | 88.97 | |
| Benchmark | Finetune Data | CHAIRi | CHAIRs | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | 80K | 72.00 | 19.70 | 32.74 | 92.27 | 9.19 |
| 158K | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 | |
| SVIT | 87.00 | 23.30 | 47.46 | 296.63 | 18.14 |
| Benchmark | Vision Encoder | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | CLIP-L-112x | 79.00 | 21.70 | 32.04 | 110.36 | 9.12 |
| CLIP-L-224x | 74.00 | 19.30 | 32.83 | 113.03 | 9.18 | |
| CLIP-L-336x | 64.00 | 16.00 | 33.37 | 108.52 | 8.92 |
| Benchmark | Vision Encoder | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | CLIP-L-224x | 79.00 | 21.70 | 32.04 | 110.36 | 9.12 |
| CLIP-L-336x | 79.00 | 18.90 | 36.00 | 111.55 | 9.19 | |
| CLIP-L-224x (SW) | 72.00 | 18.70 | 36.89 | 110.43 | 8.65 |
2.4 Vision Encoder
Intuitively, increasing image resolution enhances model’s perception of finer details, thus making the grounding of objects mentioned in the caption easier. To verify our hypothesis from the previous section, we increase the input image resolution for the vision encoder. Specifically, for our evaluation, the resolution for LLaVA7B was incremented from 224x to full resolution using a sliding window approach for efficiency, as detailed in the Appendix. Table 6 shows a constant decrease in hallucination and increase in object coverage. Additionally, we assess LLaVA with Llama 213B, varying the resolution from 112x to 336x. For the 112x112 resolution, the original image was downscaled to 112x and subsequently upscaled to 224x before being input to CLIP-Large-224x. Table 5 gives a consistent observation that larger input resolution can reduce hallucination in image captions.
Conclusion. Through our systematic analysis of object existence hallucination, we summarize several insights: (1) Enlarging the language decoder does mitigate hallucination, but improvements are not huge. (2) Expanding the volume of instruction data actually increases hallucination. Upon inspection of training data, we find certain objects described in the captions might not be grounded by the vision encoder. (3) To validate our hypothesis in (2), we show improving input image resolution significantly reduces hallucination by enhancing model grounding ability.
Reflecting on these findings, we attempt to provide an explanation for the reason of object existence hallucination in detailed captions. The process of image captioning in LMMs can be perceived as a form of information mapping or translation. Ideally, the goal is to have a direct one-to-one correspondence between objects identified in the image and those mentioned in the captions. Objects successfully grounded by the vision encoder form accurate correspondence, making this as contextual knowledge in the model, following (Neeman et al., 2023). When objects in training caption fail to ground by the vision encoder, the model learns parametric knowledge, the knowledge encoded in the model’s parameter. This kind of knowledge is the association of objects in the language with other words instead of with corresponding image object feature. During inference, when the model draws from parametric knowledge, it attempts to “guess” details not grounded by the vision module and is perceived as object existence hallucination.
3 Hallucination Controlling
Building on our prior finding that parametric knowledge can lead to hallucination, controlling or signaling the model’s use of this knowledge enables us to manage its tendency for hallucination. By incorporating a controller layer, we directly influence the output distribution linked to parametric knowledge. To this end, we developed HallE-Control, a LMM designed to control the extent of parametric knowledge within detailed captions. For this purpose, we developed two datasets to train the controller: the first captures solely contextual knowledge, while the second merges both contextual and parametric knowledge. Additionally, we can make the model signals the parametric knowledge by marking unrecognised objects in the finetuning dataset.
3.1 Data Generation
Grouping using RAM. We begin by passing MSCOCO’s ground truth objects to the open vocabulary detector RAM (Zhang et al., 2023b). RAM categorizes the objects into two groups: “grounded” (detected objects as contextual group) and “omitted” (not detected objects as parametric group). This step aims to simulate the maximum visual granularity achievable by a vision encoder in LMMs.
Contextual Data Generation. Our first dataset involves generating detailed captions using only objects from the contextual group. To do this, we feed MSCOCO source labels (including object classes, bounding boxes, and short captions) into GPT-4. We adhere to LLaVA’s caption creation pipeline and provide prompt in Appendix.
Parametric Joint Data Generation. The second dataset incorporates both contextual and parametric knowledge. Here, we begin with LLaVA’s original detailed captions and annotate objects from the parametric group with special tokens. Specifically, we enclose the “omitted” objects with brackets. Formally, if denotes the original image caption sentence and represents a set of undetected objects, our data processing can be represented as:
The purpose of bracketing the parametric objects is twofold: it serves as an indicator during inference and provides a hint during training.
3.2 Hallucination Control
Inspired by LM-Switch (Han et al., 2023), we address hallucination by adding a control parameter , e.g., for permitted imagination and for restricting imagination, as depicted in Figure 2(a). Let represent the LLM with fixed parameters: where stands for LM-head and is the output word embedding from LM-backbone. We modify , thus making output word embedding , leading to the derived model as:
The learned projector can be regarded as the transformation from a generic word space to the object sensitive word space, where word-word semantic correspondence is optimized to object correspondence, and governs the intensity of imagining related objects.
Training. To train such a controlling parameter, we leverage the contrastive training data covering both contextual and parametric datasets in Sec 3.1. For data with only contextual knowledge, we assign when inputting into the model. In contrast, for data with both contextual and parametric knowledge, we use . Notably, only the linear layer is fine-tuning throughout the training phase.
Inference. At the inference stage, can adopt any value within the interval . Specifically, an value of corresponds to minimal reliance on parametric knowledge, whereas a value of indicates a strong inclination towards such knowledge. Detailed theoretical explanation on why HallE-Control works are elaborated upon in the Appendix.
4 Experiment
| Setting | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| 158K baseline | LLaVA | 224x | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
| only ind. | LLaVA | 224x | 53.00 | 63.90 | 12.01 | – | 1.66 |
| w/o ind. | LLaVA | 224x | 57.00 | 17.10 | 37.60 | 108.94 | 7.63 |
| w/ ind. | LLaVA | 224x | 57.00 | 14.00 | 37.62 | 108.94 | 9.22 |
| 158K baseline | LLaVA Llama 2 | 336x | 64.00 | 16.00 | 33.37 | 108.52 | 8.92 |
| only ind. | LLaVA Llama 2 | 336x | 52.00 | 62.31 | 19.90 | – | 1.3 |
| w/o ind. | LLaVA Llama 2 | 336x | 52.00 | 11.62 | 34.70 | 106.94 | 7.23 |
| w/ ind. | LLaVA Llama 2 | 336x | 52.00 | 9.86 | 39.31 | 106.94 | 8.52 |
| Control value | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| 1 | LLaVA | 224x | 89.00 | 26.60 | 32.80 | 108.54 | 9.72 |
| 0.5 | LLaVA | 224x | 85.00 | 27.92 | 34.02 | 109.33 | 8.81 |
| -0.5 | LLaVA | 224x | 81.00 | 24.88 | 35.87 | 118.08 | 8.04 |
| -1 | LLaVA | 224x | 76.00 | 20.90 | 33.88 | 133.79 | 8.02 |
| 1 | LLaVA Llama 2 | 336x | 65.00 | 14.58 | 36.14 | 102.18 | 8.37 |
| 0.5 | LLaVA Llama 2 | 336x | 65.00 | 14.44 | 32.32 | 103.51 | 8.45 |
| -0.5 | LLaVA Llama 2 | 336x | 66.00 | 13.79 | 33.07 | 105.57 | 8.41 |
| -1 | LLaVA Llama 2 | 336x | 43.00 | 6.37 | 34.37 | 136.28 | 8.79 |
4.1 Finetune on Parametric Joint Data
Before we present experiments on HallE-Control, we show the upper-bound experiments on how well the model can indicate parametric knowledge. We directly finetune the LLaVA model on parametric joint data. Intuitively, the model is trained on data indicating parametric knowledge. Its output should identify parametric knowledge accurately. Specifically, the model should put a bracket around every ”guessed” objects for indication of hallucination.
Therefore, we evaluate the object hallucination in three different settings: 1. Evaluation only on indicated objects: We do CCEval only on objects inside the bracket. The result should reflect a high level of hallucination. 2. Evaluation without indicated objects: We disregard objects in bracket and calculate CCEval. The result should reflect a low level of hallucination. 3. Evaluation with indicated objects: We calculate CCEval on all objects. Due to modification of the settings, we slightly change definition of CHAIR scores in CCEval as detailed in Appendix.
Evaluation only on indicated objects. The hallucination level for indicated objects, , is 63.90 for LLaVA7B and 62.31 for LLaVA13B. It is considerably higher than the baselines all other models. Concurrently, their coverage is 12.01 and 19.09 for LLaVA7B and LLaVA13B, respectively, which both of them significantly lower than the coverage of 33.58 for LLaVA7B baseline model. The experiment results show the object within special tokens has significantly higher hallucination rate and lower coverage rate which support our assumption that the object inside the indication tokens are objects from parametric knowledge.
Evaluation without indicated objects. For objects outside of the special token scope, we find hallucination is markedly reduced. CHAIRs decreased from 82 to 57 which is 30.5% percent improvement; CHAIRi decrease 32.4%, from 25.3 to 17.10, compared to the baseline. This suggests that the model is less prone to make erroneous assumptions for objects not marked by brackets. This is interesting because the model perfectly captures the intention of marking parametric objects in training data and replicate the behavior during inference.
Evaluation with indicated objects.: We observe a significant decline in the hallucination without any reduce in object coverage. LLaVA7B CHAIRi improved from 25.30 to 14.00 which has 44.66% improvements. For LLaVA13B CHAIRi improved from 16 to 9.86 which also has 38.38% improvements.
4.2 Hallucination Controlling
During model inference, we select 4 different , ranging from to . As shown in Table 8, we evaluate HallE-Control7B and HallE-Control13B model, which use LLaVA as backbone. For the 7B model, we train it by removing the special token in parametric joint data, showing that indication is not a necessary condition for control module to work. means the control module trained purely on contextual only data, which try to minimize the hallucination, where maximize parametric knowledge output. The results show that as increase, CHAIRi increase from 20.90 to 26.6, the coverage keeps at a similar level.
For the 13B model, we keep the indication inside the parametric joint data. The HallE-Control13B achieves the best in object existence hallucination metric. With control value set to -1 and indication, we have versus baseline’s 64 and vs baseline’s 16, and the coverage of the model is not decreasing.
5 Related Work
Large Multimodal Models (LMMs).
The rapid advancements in Large Language Models (LLMs) (Touvron et al., 2023a; Chung et al., 2022; Touvron et al., 2023b; Anil et al., 2023; Driess et al., 2023; Scao et al., 2022; OpenAI, 2023) combined with a surge in open-source initiatives, have paved the way for the emergence of extensive vision-language models (Liu et al., 2023d; Goyal et al., 2017; Zhu et al., 2023; Sun et al., 2023; Ye et al., 2023; Bai et al., 2023; Chen et al., 2023; Peng et al., 2023). LLaVA introduces the concept of integrating a simple projector during LLM fine-tuning. Chatspot (Zhao et al., 2023b) follow LLaVA’s model structure, but embed region of interest into instruction data. GPT4RoI (Zhang et al., 2023a) and Shikra (Chen et al., 2023) add grounding tasks to LLaVA structure models. Instead of using detector to provide region information to the model, we use detector to filter objects for alignment between vision and language information. Additionally, BLIP2 (Li et al., 2023d) and InstructBLIP (Dai et al., 2023) present Q-former-based LMMs. Multimodal-GPT (Gong et al., 2023) and Otter (Li et al., 2023b) aims to improve OpenFlamingo’s (Alayrac et al., ; Awadalla et al., 2023) directive adherence. mPLUG-Owl (Ye et al., 2023) suggests a two-step method: first train vision models, then refining the language model using techniques like LoRA. Our work utilize a linear layer to control object existence hallucination within LMMs.
Evaluation on LMMs.
The evaluation of large vision-and-language models (LMMs) (Yu et al., 2023; Liu et al., 2023b; c) is challenging due to the intricate nature of generation tasks they undertake. Some of the VQA-based benchmarks (Antol et al., 2015; Hudson & Manning, 2019; Gurari et al., 2018) require models to identify objects, colors, or quantities, while others (Liu et al., 2023e; Li et al., 2023c; Lu et al., 2022) offer multiple-choice questions. POPE (Li et al., 2023e) and MME (Fu et al., 2023) evaluate object hallucination using paired yes/no questions on existence, color, counting, OCR, and etc. While VQA-based benchmarks are cheap and straightforward, we find them cannot accurately reflect object hallucination for detailed captions. Besides VQA benchmarks, ROUGE (Lin, 2004; Elliott & Keller, 2014) use n-gram to evaluate similarity between ground truth and model inferences. CIDr (Vedantam et al., 2015) is a triplet-based method of collecting human annotations to measure consensus. CHAIR (Rohrbach et al., 2018) evaluates caption hallucination with object concepts. These methods are constraint by ground truth length or word variance and cannot clearly reflect hallucination with object coverage information. Wang et al. (2023) uses a language model to predict whether the caption exist hallucination, which is cheaper than GPT-4. Our work introduces CCEval, including CHAIR metrics, object coverage, average sentence length and number of objects to overcome limitations of previous evaluations.
Hallucination
Hallucinations (Ji et al., 2023a; Shi et al., 2023; Lin et al., 2021) have been widely studied in traditional NLG (Ji et al., 2023b) tasks, including machine translation (Zhou et al., 2020; Lee et al., 2019), data-to-text (Rebuffel et al., 2021; Kasner & Dušek, 2022; Lee et al., 2022), summarization (Cao et al., 2022), dialogue (Dziri et al., 2022) and QA (Shuster et al., 2021). For LMMs, previous studies have been mainly focusing on object hallucination (Marino et al., 2019; MacLeod et al., 2017; Li et al., 2023a; e). POPE (Li et al., 2023e) reveals object existence hallucination may related with label distributions, such as object co-occurance. Earlier than POPE, Biten et al. (2022) balance object co-occurance to decrease hallucination. LRV (Liu et al., 2023a) finds the cause of hallucination in VQA benchmarks, especially unbalanced answer distribution and lack of negation information. Our work raise another important cause: misalignment between the vision and language information captured by models. More interestingly, we can explain balancing labels in Biten et al. (2022) as trying to weaken the parametric knowledge caused by the misalignment.
6 Conclusion
In summary, this study delves deep into the object hallucination phenomena within the detailed captions of LMMs, advancing understanding of the accuracy and unwarranted inference in describing visual details. We introduce a novel and comprehensive evaluation method for object existence hallucination in detailed captions. We conduct an in-depth and component-wise analysis of LMMs, meticulously examining each element that might result in hallucination. We further identify an alignment issue between the vision encoder and the instruction data. To alleviate such hallucination, we introduce controlling parameters over LMMs to condition the inference of objects.
Appendix A Appendix
A.1 Why not all Hallucination Bad?
In our study, we define ”object existence hallucination” to be a phenomenon where a description makes reference to objects that is not present in the image. However, such hallucinations, when properly harnessed, can be regarded as instances of imagination. Human beings frequently use imagination to successfully accomplish tasks, often without even realizing it. Here, we present several scenarios in Table 9, Table 10, and Table 11 in which imagination and judicious inference prove to be beneficial for various downstream applications, including robotic manipulation and content moderation, among others. Our work successfully exercise control over the extent of generalization rather than attempting to eliminate all instances of imagined objects. More importantly, our model can indicate imagined objects with [object] marker.
A.2 More Details on Analyzed Models
In Section 2, we conduct analysis on different LMMs. Here, we provide more details of each model:
LLaVA uses a linear projector to map visual token as a soft-prompt into LLM input tokens. LLaVA has a two-stage training, where the initial stage focuses on simple caption pretraining solely for the linear projector, while the subsequent stage finetunes both the projector and LLM on instruction data. Instruction data leverages language-only GPT-4 by inputting visual ground truth from COCO dataset.
InstructBLIP adopts the BLIP-2 architecture, and is distinguished by its training of a Q-former, which bridges the frozen vision encoder and LLM. InstructBLIP’s instruction fine-tuning spans across 26 distinct datasets.
Shikra mirrors LLaVA’s model structure. It eliminates the pretrain stage, but introduce grounding task during finetuning. Shikra is trained on multiple datasets like InstructBLIP.
A.3 Training Data Quality
We sampled three images from the MSCOCO dataset, as illustrated in Table 12. For each image, we present the visual content, a detailed caption generated by GPT-4 based on bounding boxes and regional captions, and the object ground truth labels derived from the MSCOCO dataset annotations.
First Image. This image showcases three remote controls, posing a unique challenge. The ambiguity lies in distinguishing the type of remote, be it for video games or televisions. Additionally, the remotes near the wall are relatively small, making them harder to see. A person is partially visible in the image, with only their knees being evident. There is also an incomplete bottle on the image’s left side.
Second Image. A significant issue with this image is the individuals visible behind a window. Not only detection models struggle to recognize them, even for human observers, counting the individuals inside the train is challenging. This particular issue is prevalent in many of MSCOCO’s traffic-related images.
Third Image. This image depicts a table around which two individuals are seated. Close observation is required to recognize both individuals. A clear indication of one person is a pair of hands, while the other individual is considerably harder to spot. Additionally, there seems to be an annotation error in the ground truth labels: it indicates only one bowl, neglecting to include plates and other items. Upon closer inspection, there are four plates.
| Image | Caption | GT Labels |
The image depicts a cozy entertainment room featuring a flat screen TV situated against the wall. A dining table can be seen next to the TV, with some bottles and a cup placed on top of it. Multiple video game controllers, including Wii remotes, are also present on the dining table, suggesting that people are using the TV to play video games. To the far right of the scene, another remote is located close to the wall. A person appears to be sitting or standing behind the dining table, likely either watching the game or waiting for their turn to play. The assortment of bottles on the table suggests that the guests are enjoying drinks during their video game session. | tv: 1 cup: 1 remote: 3 bottle: 4 person: 1 dining table: 1 |
In the image, an orange mass transit trolley is making its way through a city. A person is crossing the street in front of the trolley while holding a garbage bag, appearing to be cautious about the approaching vehicle. Another person is standing close to the person crossing the street, and there are two more individuals nearby. In the scene, various vehicles surround the trolley, including cars, buses, and trucks. One of the cars is parked right behind the trolley, while another is situated farther back. Two buses can be seen, with one staying behind the trolley and the other on the right side of it. A truck is also present at the far left side of the scene. A woman nearby is holding a handbag, completing the busy urban setting. | car: 3 bus: 2 truck: 1 train: 1 person: 5 handbag: 1 |
The image shows a table in a restaurant, which appears to have been recently used for a meal. The table is set with four place settings, including white dishes. On the table, there are multiple cups, a bowl, and a mix of silverware like forks and spoons laid out. Some of the dishes look dirty, with used napkins and eating utensils scattered around. In the background, two people can be seen, one situated on the left side and another on the right side behind the table. There are also chairs located near the table, with one chair positioned close to the left side and another chair closer to the right side of the frame. | cup: 5 spoon: 3 fork: 2 person: 2 chair: 2 bowl: 1 dining table: 1 |
A.4 CCEval
In this paper, we introduce CCEval, a GPT-4 assisted Caption Concept Matching Evaluation. CCEval provides concept matching method and metric that can be extended to larger scale. Here, we detail the process of evaluation:
Step 0. Collect ground truth objects from annotations in a list.
Step 1. Run the model through the evaluation set of 100 images from Visual Genome. Specifically, prompt the model with ”Describe the image in detail” and obtain 100 captions as model outputs.
Step 2. Extract objects mentioned in 100 captions using GPT-4.
Step 3. For each caption, match the extracted objects with ground truth objects using GPT-4.
GPT-4 prompts are provided in the next section. After Step 3, we get the following information: matched objects (no hallucination), unmatched objects (hallucination), total objects mentioned, total ground truth objects. Our metric calculates:
Note that we do not strictly enforce consistent constraints (same coverage, average length, and average objects) across models, neither this is realistic to achieve. Although our intention is to make the model comparable and push the model to express more objects, it is sometimes not possible for some of the models, such as BLIP2, to express detailed captions, and our coverage, average length, and average number of object metrics in CCEval can accurately reflect if any model has such limitations.
In the main text, ”consistent constraints” are mainly achieved through prompting. We use the same prompt - “describe this image in detail” - across all four models in Table 2. Simply prompting can result in similar length and number of objects captions in this case since the instruction finetuning data for detailed captioning all comes from a similar source (LLaVA synthetic data). For models with different instruction finetuning data, we suggest further adjustment of the prompt or set the min_length and max_length in the inference function.
A.5 Prompt
Table 13 is the prompt we use for caption object extraction.
Table 14 is the prompt we use for hallucination object matching.
Table 15 is the prompt we use for finding ground truth object coverage.
A.6 Sliding Window
We pad the original image to dimensions of . Then, we divide the image into a grid, where each cell measures . The encoder processes these cells sequentially, starting from the top-left and moving towards the bottom-right. The visual tokens from each cell are then concatenated.
A.7 Experiment Settings
The LLaVA we used in all experiments is pretrained on on LCS-558k which is subset of LAION (Schuhmann et al., 2021), CC (Sharma et al., 2018) and SBU (Ordonez et al., 2011) data, and finetuned on Instruct-158K instruction data. We use Vicuna version 1.3 as initialized language decoder and CLIP-Large as vision encoder. The 158K finetuning data consists of detailed caption, complex reasoning, and conversation.
We employed the RAM detector, specifically the RAM-14M variant, which uses a Swin-Large backbone. In the data generation stage, we focused on the ’detailed caption 23K’ file from the LLaVA Instruction set comprising 158K entries. This file was generated using a specific prompt provided by LLaVA repository111https://github.com/haotian-liu/LLaVA/tree/main/playground/data/prompts/detail_description. Our preprocessing involved adding brackets ’[]’ around objects filtered by the RAM detector to uniquely identify them in the captions.
Hallucination control only involves only finetuning a linear layer added before layer. We freeze all other layers and only finetune control layer with 3 epochs. The dataset is generated data with associated value. We have 10K contextual only detailed caption data and 23K parametric joint detailed caption data.
A.8 More Experiments on LLaVA 1.5
| Setting | LLM | Resolution | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| 665K baseline | LLaVA | 336x | 66.00 | 17.84 | 34.43 | 103.63 | 8.35 |
| 665K baseline | LLaVA | 336x (SW) | 64.00 | 15.54 | 34.39 | 101.33 | 8.30 |
| w/o ind. | LLaVA | 336x | 54.00 | 11.24 | 34.00 | 103.62 | 6.76 |
| w/ ind. | LLaVA | 336x | 54.00 | 8.91 | 35.94 | 103.62 | 8.53 |
The experiments on the main text are based on LLaVA 1.3, we provide experiments on LLaVA 1.5 in Table 16. The results are consistent with the conclusion in the main text and hallucination is less in this new version.
A.9 Modified CCEval of Section 4.1
1. Evaluation only on indicated objects
2. Evaluation without indicated objects
3. Evaluation with indicated objects
A.10 Qualitative Results of HallE-Control
This section shows qualitative results for HallE-Control in Table 17, Table 18, Table 19, Table 20, and Table 21. means no imagination and means max imagination. Both models can output indication.
A.11 Theoretical Explanations
HallE-Control Formulation
HallE-Control follows LLaVA’s training strategy, which freezes vision encoder and language model, only finetune the projector for the first stage. The visual encoder transfer image to visual features:
The projector connects image feature to the word embedding space:
In the second stage, both projector and language model is trained. The model is trained on two type of data: 1. Contextual only data. 2. Contextual and parametric combined data.
Language model trained on contextual only data has a distribution initiating from . The other language model trained on combined data has a distribution initiating from
With the Theorem 1 in LM-Switch (Han et al., 2023), under the same assumptions, there exists an matrix , transforming a word embedding to , which is equivalent to let a LM simulate the text distribution initiating from another distribution.
Inspired by the Theorem, we propose a linear transform in word embedding space for LMM. Let be the finetuned LLM, with control layer/projector denoted as , we replace each word embedding with , making the new language model as HallE-Control’s language decoder. is adjustable from -1 to 1. We assign and fit contextual only data; assign with the contextual and parametric combined data. After finetuning the HallE-Control, the user only needs to specify a control value and do normal vision language task like image captions. We use maximal likelihood as the training objective.
Continuous Control
The design of LM-Switch maintains a linearity guarantee, with proof of the control model’s distribution is close to a linear interpolation. Let be the maximum eigen-value of W. When varying s value,
distribution of the control model is close a linear interpolation of and , meaning that the model distribution changes linearly. Therefore, our method can take any between 1 and -1.
For HallE-Control, the core idea is: Assuming LLM is good enough to represent an equivalent distribution with HMM; there exist an matrix W, so that after transferring word embedding to , the LLM’s originally simulate the text distribution starting with initial state will turn to be equivalent to a distribution starting with initial state .
According to experiments in Section 4.1, it reveals language models can distinguish inference object and observed objects by adding special tokens. Because, decoder based language models generate sequences in an auto-regressive way. We followed the proof in Han et al. (2023): if we assume O as our observation space which means a set of m observations. We assume as a contextual vector, as word embedding. We can represent word logits as E. In attention mechanic, it will pass through a softmax operator to get the distribution over words, We use the same assumption in LM-Switch method to assume a linear formulation and let the conditional probability in language model . We can get the full probability by using chain-rule: . We assume our LLMs are good enough to represent an equivalent distribution with HMM and full column-rank for E, p(o).
Our conclusion is: Applying a linear transformation on word embedding space is equivalent to a shift from one initial condition to another. This is the reason that we want to shift a language model with distribution produce higher inference imagination conditional probability to a lower one.
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000146988.jpg)
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000157923.jpg)
![[Uncaptioned image]](extracted/2310.01779v3/figures/000000192920.jpg)